OpenStudy (anonymous):

Mike observed that 75% of the students of a school liked skating. If 35 students of the school did not like skating, the number of students who liked skating was ______. (only put numeric values, no other symbols)

OpenStudy (anonymous):

IF 75% liked skating, what percent did not like skating?

OpenStudy (anonymous):

check mines 35 = 0.25x as percentages are decimals out of 1 Solve by dividing 35/0.25.. x=140 So, 140-35=105.??? thats the answer im just makin sure

OpenStudy (anonymous):

105

OpenStudy (anonymous):

Looks good to me :) Well done

OpenStudy (anonymous):

thats rite 0_0

OpenStudy (anonymous):

Yep, that's correct. There's a few ways to do this problem, but that's the most general (which means it will work for other percentages).

OpenStudy (anonymous):

o ty ! can you help me with another one =D

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Solve for H: A = L • W • H A)H = A times W over L B)H = W times L over A C)H = L times A over W D)H = A over quantity L times W

OpenStudy (anonymous):

Have you tried this one, as well?

OpenStudy (anonymous):

\[A = LWH\]\[\frac{A}{L} = \frac{LWH}{L}\]\[\frac{A}{LW} = \frac{LWH}{LW}\] \[H = \frac{A}{LW}\]

OpenStudy (anonymous):

ty :]

OpenStudy (anonymous):

My pleasure :)